OpenStudy (anonymous):

For the following geometric sequences, determine a1, r, a5, and a6. 8,4,2,1,1/2,...

OpenStudy (zzr0ck3r):

do you know what a1 is?

OpenStudy (anonymous):

8?

OpenStudy (zzr0ck3r):

correct. what is r?

OpenStudy (anonymous):

0.5 or 1/2 im not sure if it matters or not.

OpenStudy (zzr0ck3r):

they are the same thing, but we normally use fractions unless asked otherwise

OpenStudy (zzr0ck3r):

and a5?

OpenStudy (anonymous):

1/2?

OpenStudy (zzr0ck3r):

so a2=1/2*a*1 a3=1/2*a2 what is a6?

OpenStudy (anonymous):

1/4??

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

\(\huge \checkmark\) and you are done

OpenStudy (anonymous):

okay, thank you but can you help me with a couple of examples please?

OpenStudy (anonymous):

1,10,100, 1,000, 10,000, ...

OpenStudy (anonymous):

a1=1 r= 10?

OpenStudy (anonymous):

hello? :(

OpenStudy (anonymous):

Pretty \(\color{green}{\checkmark}\)..

OpenStudy (zzr0ck3r):

correct @Jmoya_04 maybe don't use the commas to write 1000 :)